OpenStudy (yamyam70):

solve the exponential equation , (3)^x +2 = 7 log 3 ^ x +2 = log 7 x + 2 log 3 = log 7 whats next ?

OpenStudy (anonymous):

divide both sides by log3

OpenStudy (yamyam70):

x + 2 = log 7 / log 3

OpenStudy (anonymous):

now subtract 2 from both sides

OpenStudy (yamyam70):

x = (log 7/ log 3 ) -2 correct?

OpenStudy (yttrium):

you can do like this 3^x + 2 = 7 3^x = 5 x = log(5)/log(3)

OpenStudy (anonymous):

yes its correct. its the value of x. @yamyam70

OpenStudy (yamyam70):

-0.2 ,

OpenStudy (yamyam70):

thanks much @rizwan_uet :))

OpenStudy (anonymous):

welcome